I need help solving this word problem...I have been trying for 2 hours.
In 1920, the record for a certain race was 45.4 sec. In 1970, it was 44.4 sec.
Let R(t)=the record in the race and t=the number of years since 1920.
Find a linear function that fits the data
R(t)=
Round to nearest hundredth
What is the predicted record for 2003____sec
What is the predicted record for 2006?
In what year will the predicted record be 43.50 seconds?
Thanks so much!

Answer by mananth(16946)

year Record time
1920 45.4 seconds
1970 44.4 seconds
Difference
50 -1
rate of reduction per year =-1 /50 = -0.02

R(t)= 45.4-0.020t

2003 t=83

R(t)= 45.4 + -0.020 * 83

R (83)= 43.74 seconds
------------

R(t)= 45.4-0.02t
43.5=45.4-0.02t
45.4-43.5=0.02t
1.9=0.02t
t= 1.9/0.02
t=95 years
1920 +95 = 2015
year 2015 it will bve 43.5 seconds